OpenStudy (anonymous):

Where does the graph of y = 6x2 − 11x − 10 cross the x-axis Answers must be fractions? Dont get it

OpenStudy (anonymous):

........

OpenStudy (anonymous):

...............and...........

OpenStudy (anonymous):

At the point, where the graph crosses the x-axis, the value of the y-coordinate is 0.. therefore to find the value of x at that point (where curve crosses the x-axis), your equation now becomes \[0 = 6x ^{2}-11x-10\]

OpenStudy (anonymous):

I know i just dont know how to get answers in fraction form

OpenStudy (campbell_st):

well the easiest way is to use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a =6, b = -11 and c = -10 substitute them and evaluate

OpenStudy (radar):

Factor it: (3x+2)(2x-5)=0 You gonna get fractiolns when you solve it.

OpenStudy (anonymous):

i did..... i need answers in fraction form as I stated multiple times

OpenStudy (anonymous):

if 3x+2 = 0, you will get a fraction if you solve for x... same with the other factor that radar gave you

OpenStudy (anonymous):

use the quadratic formula above, or if it factorizes easily then factor it out. Solve for x. You should get two "fractions".

OpenStudy (anonymous):

How did you get that @radar

OpenStudy (anonymous):

If you solving for the X intercept. It means Y is going to be 0. So set y to 0. Then you get =0. Solve for x using two methods: Factorizing like radar did, or plugging the values from the quadratic into the QUADRATIC FORMULA to get your two x values.

OpenStudy (anonymous):

@adziz there really is no need for your smartassery

OpenStudy (radar):

I did it by looking at 6 breaking it down to factors 3 and 2, then doing some minor trial and error by multiply ing each of those by the factors of 10 which are 5 and 2 and noting the signs, come up with a product that when summed equals -11. Takes practice but it becomes pretty reasonable.

OpenStudy (anonymous):

ok